# OpenFake: An Open Dataset and Platform Toward Large-Scale Deepfake Detection

### v1

@@ 3.1 Social media content
As part of a subsequent study, we have collected social media posts—including images and videos—using a cloud-based, real-time multimedia data collection pipeline deployed on AWS. This pipeline continuously monitors and filters public content across three platforms: X, Reddit, and Bluesky. The data capture spans the period leading up to and including the 2025 Canadian federal election, enabling analysis of online discourse during this critical time.
Systematic manual verification of 650 randomly sampled images collected during a 24-hour period identified 51 distinct deepfake instances (7.8%). Verification methodology involved source authentication and reverse image search techniques. This detection rate should be considered conservative, as the actual prevalence of deepfakes within the dataset is likely higher due to limitations in manual detection methods. Notably, many of these deepfakes prominently featured political figures. For instance, fabricated scenarios involving leading candidates significantly distorted their public perception. Even when certain deepfakes were clearly artificial (Figure˜2 a), featuring visual inconsistencies or unrealistic scenarios, they effectively reinforced existing biases or narratives, resonating more profoundly with audiences compared to textual misinformation alone (Ecker et al., 2022; Hameleers et al., 2020; Vaccari and Chadwick, 2020). When the deepfake is convincing, the potential for harm increases substantially. Viewers are more likely to accept the content as genuine, especially when the image aligns with their preexisting beliefs or emotional biases. For example, as shown in Figure˜2c, a photorealistic yet entirely fabricated image of a political figure in an implied context, despite lacking any overt manipulation cues, can mislead viewers through association alone.
Beyond fabricated portraits of politicians, misinformation via images also extended to signs, banners, and manipulated visuals of significant news events such as protests or catastrophes as seen in Figure˜2 d) and e). These images often appeared alongside carefully crafted textual content designed to reinforce misleading narratives as exemplified in Figure˜2 b). Advanced Vision-Language Models (VLMs) enable the extraction of any textual content present in images, so our dataset does not need to explicitly target images with visible text. Any such content can still be leveraged for multimodal misinformation detection. Prior work has shown that combining VLM-extracted text with Large Language Model (LLM)-based fact-checkers can enhance automated verification systems (Tahmasebi et al., 2024; Raza et al., 2025; Li et al., 2024).
Altogether, our analysis highlights that deepfakes, while often focused on portraits, extend far beyond facial imagery. They frequently depict fabricated news events, disaster scenes, protest visuals, and manipulated political symbols—forms of visual misinformation that can be equally, if not more, influential. These findings directly inform the design of our dataset: to support generalizable detection, a deepfake benchmark must move beyond faces to capture the full breadth of misleading visual content. This motivates the use of a targeted filtering pipeline and underscores the need for robust, scalable detection systems equipped to handle the multimodal nature of modern misinformation.
@@ 3.2 Human Perception Study
To assess the difficulty of detecting deepfakes generated by different models, we conducted a simple human study. A total of 100 participants completed the survey, each viewing a randomized set of 24 images. The set consisted of 12 real photographs and 12 synthetic images—2 from each of the six generation models: GPT Image 1 (OpenAI) (OpenAI, 2025), Imagen 3 (Google) (Baldridge et al., 2024), Grok 2 (xAI), Flux.1.0-dev (Black Forest Labs) (Labs, 2024), Stable Diffusion 2.1 Rombach et al. (2022), and Stable Diffusion 3.5. All synthetic images were generated from the same prompts as their real counterparts, using the text automatically extracted by our pipeline described in Section˜4. We did not hand-pick or optimize the generations, aside from removing names and content restrictions when necessary. Each prompt was only shown once to a given participant. This ensured that responses reflected a fair and diverse exposure across the dataset. In total, the survey contained 168 unique images. The survey description, code, and images are available at https://github.com/vicliv/deepfake-survey.
Model
@@
Flux.1.0-dev 2024 Open-source 0.689 0.562 0.792 0.812 0.917 1.000 1.000
SDv3.5 2024 Open-source 0.709 0.521 1.000 0.938 0.750 1.000 1.000
Grok 2 2024 Proprietary 0.811 0.771 1.000 0.938 0.208 1.000 0.875
SDv2.1 2022 Open-source 0.879 0.646 1.000 0.938 0.917 1.000 1.000
0.714 0.524 0.417 0.427 0.804 0.750 0.990
@@
Table 2: Survey results showing human true positive rate (TPR) for synthetic images and true negative rate (TNR) for real images. A score near 0.50 indicates chance-level performance. SwinV2 trained on our data achieves high accuracy on detecting fakes from the models it was trained on, but underperforms on unseen or low-quality deepfakes. Grok 2 is kept out of the training data to verify SwinV2’s out-of-distribution performance (only for the survey results).
@@
The results in Table˜2 reveal several key insights into the current landscape of synthetic image realism and human perception. Most notably, Imagen 3 from Google achieved the lowest human accuracy (48.5%), performing no better than random guessing. GPT Image 1, OpenAI’s recent image generation model, also ranked among the most deceptive, with nearly one-third of its images going undetected. By contrast, Stable Diffusion 2.1 (SDv2.1), which is widely used in current benchmarks, had the highest human detection rate (87.9%) and was easily flagged by all detectors. This suggests that while open-source models are still relatively easy for both humans and machines to spot, advanced proprietary models exhibit an exceptionally high degree of realism and consistency, underscoring the importance of including them in benchmarks when evaluating the limits of human and machine perception. This demonstrates that proprietary models are setting a new bar in generation quality, one that current deepfake detection benchmarks, largely trained on older or open-source models, fail to represent.
@@
However, this perceptual difficulty is not limited to humans. Baseline deepfake detectors also failed to consistently identify these advanced fakes. The CLIP-D-10k+ method (Cozzolino et al., 2024), which fine-tunes a linear classifier on CLIP embeddings, performed close to random on several models (e.g., 45.8% on Imagen 3, and GPT Image 1), and failed to distinguish real images entirely (47.9% TNR). Similarly, the Corvi2023 method (Corvi et al., 2023), which uses curated handcrafted features for detection, fared better on some open-source models (e.g., 100% on SDv3.5), but completely failed on real images (0.0% TNR) and newer proprietary content like GPT Image 1 (50.0%). In contrast, our SwinV2 baseline, trained directly on a curated mix of real and synthetic images from models included in our dataset, achieved 99.0% overall accuracy and perfect performance on its in-distribution models while having a reasonable score of 87.5% on Grok 2 (kept out of the training for this test). While this result demonstrates that deep networks can learn to detect even the most realistic fakes with sufficient supervision, it also highlights the brittleness of existing methods when faced with novel or unseen generative sources. The two additional SwinV2 baselines confirm this pattern: models trained on smaller or differently distributed data (GenImage and Semi-Truths) showed weaker generalization to modern deepfakes, with accuracy dropping to 80.4% and 75.0%, respectively. These results reinforce the importance of training on a large, diverse dataset like OpenFake to ensure robustness against emerging generation techniques.
In summary, both humans and the older detection systems we tested struggle to identify deepfakes generated by modern diffusion and transformer-based models. These findings call into question the utility of outdated benchmarks and reinforce the need for detection datasets that reflect the state of the art. Our study motivates a shift toward more diverse, high-fidelity, and license-aware synthetic corpora, paired with regular human benchmarking, to ensure that future detectors remain robust in the face of rapid generative progress.
@@ 4 Dataset Overview & Collection
Figure 4: t-SNE visualization of CLIP vision embeddings for 3,000 test images, including both real and synthetic images from a few generative models. Each point corresponds to an individual image, and colours indicate the generative model (or “real” for authentic images).
OpenFake combines a large repository of real images with a diverse collection of high-quality synthetic counterparts generated by multiple state-of-the-art models. In Appendix A, Table˜5 presents in details the key statistics of the dataset. Figure˜4 offers a qualitative view of the underlying distribution of some of the images. The substantial overlap between real and synthetic samples in the CLIP feature space highlights their semantic alignment, suggesting that synthetic images effectively mimic the distribution of real-world content. This shows that the prompt generation pipeline is working as intended. An overview of the data collection and generation process is provided in Figure˜1.
@@
We extract metadata from the LAION-400M dataset (Schuhmann et al., 2021), which we selected due to its broad representation of internet-sourced images—the same domain where visual misinformation typically circulates. Additionally, this dataset was likely included in the training data of the text-to-image models used to generate the synthetic images, which should theoretically make it more difficult for detectors to distinguish between real and fake images. More importantly, these images preserve real-world compression artifacts, which are crucial for training detectors that operate in the wild. While LAION may contain some synthetic images, we expect this contamination to be minimal, as the dataset primarily consists of content from 2014–2021, before the public release of diffusion models in 2022. After scraping LAION, we filter image–caption pairs using a vision–language model (Qwen2.5-VL (Bai et al., 2025)). As described by the prompt used to query the model in Appendix˜B, an image is retained if it is identified as depicting either (i) real human faces or (ii) politically salient or newsworthy events. For every retained image, we generate a more detailed caption to use as prompt input for text-to-image models. These $3\,\mathrm{M}$ prompts are also publicly released and form the basis of the prompts shown to users of our crowdsourcing platform OpenFake Arena.
We filtered real images using Qwen2.5-VL, selected for its trade-off between speed and quality. To prevent detection shortcuts, we excluded LAION images smaller than $\times512$ pixels from the released train/test sets, as lower resolutions introduced compression artifacts that made detection artificially easier. Full prompts used for filtering and captioning are provided in Appendix˜B, and additional details on generation and compute resources are in Appendix˜D.
@@
We generated images from a diverse set of state-of-the-art generators using samples from our prompt bank: Stable Diffusion 1.5/2.1/XL/3.5 (Rombach et al., 2022), Flux 1.0-dev/1.1-Pro/Schnell (Labs, 2024), Midjourney v6/v7 (Midjourney, 2024, 2025), DALL·E 3 (OpenAI, 2023), Imagen 3/4 (Baldridge et al., 2024; Google Cloud, 2025), GPT Image 1 (OpenAI, 2025), Ideogram 3.0 (Ideogram AI, 2025), Grok-2 (xAI, 2024), HiDream-I1 (HiDream-ai, 2025b, a), Recraft v3 (Recraft, 2024), Chroma (Lodestones, 2025), and 10 community variants (Finetuned or LoRA) of Stable Diffusion 1.5/XL and Flux-dev. All images are produced at $\sim$1 MP resolution with varied aspect ratios (9:16, 16:9, 1:1, 2:3, 3:4, etc.), mirroring common social-media formats. Because several proprietary sources impose “non-compete” clauses, those subsets are released under a non-commercial license.
@@
We construct the train/test split by sampling 1,000 images per generative model (with the exception of out-of-distribution models, which contribute between 200 and 600 images) along with the corresponding number of real images, yielding a test set of roughly 60,000 images (about 3% of the dataset assuming balanced classes). The remaining images are allocated to the training set. To ensure balance, each model is equally represented in the test split, and real images are matched accordingly. The rest of the real images and prompts are provided in a CSV file, with the real images accessible through their URLs. As OpenFake Arena expands and more synthetic images are collected, additional real images will be incorporated to preserve parity between real and synthetic domains in the train/test splits. All assets are hosted on the HuggingFace Hub at ComplexDataLab/OpenFake. All images have their associated prompts and model name as metadata, which can be used for model attribution.
Limitations.
Our dataset reflects limitations from its sources. Real images from LAION are skewed toward Western, pre-pandemic content, and generative models carry aesthetic and cultural biases. We mitigate risks via filtering, Arena submission review, and safeguards against misuse, but adversarial content and prompt noise remain challenges. More details are in Appendix˜C.
@@ 5 Crowdsourced Adversarial Platform
The Arena features a real-time leaderboard to gamify the experience and incentivize participation. The detector is periodically retrained with newly collected data, enabling continual improvement. Submitted images are periodically reviewed. This human-in-the-loop setup transforms model drift from a challenge into a feature, allowing the benchmark to evolve organically alongside the state of generative models. Implementation details and screenshots of the arena are in Appendix˜E.
@@ 6 Baseline Detector Benchmarks
We benchmark a broad set of representative detection paradigms, spanning supervised, semi-supervised, and zero-shot approaches. Our primary supervised detector is SwinV2-Small (Liu et al., 2021b), a vision transformer pretrained on ImageNet-1k (Deng et al., 2009) and finetuned on our dataset. We further evaluate two SwinV2 variants trained on external datasets (GenImage and Semi-Truths), a ConvNeXt baseline trained on DRCT taken from (Chen et al., 2024a), and an EfficientNet-B4 baseline trained on FaceForensics++ from (Yan et al., 2023). For semi-supervised detection, we include the CLIP-Based Synthetic Image Detector (Cozzolino et al., 2024), which applies a linear probe over CLIP embeddings with minimal training data. For zero-shot detection, we use InternVL (Chen et al., 2024c) directly without finetuning. [^4] We use the InternVL3-38B; Code for all scripts and baselines https://github.com/vicliv/OpenFake Finally, we also test the handcrafted detector of Corvi et al. (2023) and a hybrid fusion baseline that averages predictions from CLIP and Corvi2023. Together, these baselines cover a diverse range of architectures, training regimes, and prior benchmarks, allowing us to compare detectors trained on OpenFake against both legacy and contemporary approaches.
Results on OpenFake.
As shown in Table˜3, the SwinV2 trained on OpenFake is near-perfect, confirming that modern classifiers are highly reliable when trained on the evaluation distribution. The SwinV2 trained on GenImage is the next strongest and handles many shared open-source families, but degrades on newer proprietary generators (Grok 2, Midjourney 6, Flux-1.1 pro, DALL·E 3), reflecting a distribution gap in the fake images. The SwinV2 trained on Semi-Truths achieves high TPRs yet misclassifies many real images, which is consistent with its training data focused on edits rather than full generation. Legacy or narrow baselines (ConvNeXt/DRCT, EffNet-B4/FF++) underperform, CLIP-D-10k+ is middling, and Corvi2023 largely predicts “real,” while zero-shot InternVL is better than older baselines but still trails supervised models. Overall, robust performance requires training on the correct, broad, and up-to-date image distribution.
Metric / Model SwinV2 (Ours) SwinV2 (GenImage) SwinV2 (Semi-Truths) ConvNeXt (DRCT) EffNet-B4 (FF++) CLIP- D-10k+ Corvi 2023 InternVL-3 (zero-shot)
Training Dataset OpenFake GenImage (Zhu et al., 2023) Semi-Truths (Pal et al., 2024) DRCT (Chen et al., 2024a) FF++ (Rössler et al., 2019) Custom (Cozzolino et al., 2024) DMD (Corvi et al., 2023)
Real Images TNR 0.995 0.955 0.689 0.777 0.516 0.703 0.998 0.431
Stable Diffusion 1.5 1.000 0.936 1.000 0.447 0.529 0.579 0.000 0.849
Stable Diffusion 2.1 1.000 0.998 0.999 0.482 0.453 0.717 0.011 0.900
Stable Diffusion XL 1.000 0.956 1.000 0.426 0.507 0.438 0.001 0.814
Stable Diffusion 3.5 1.000 0.982 1.000 0.324 0.466 0.406 0.000 0.796
@@
Test set Metric Train OpenFake Train GenImage Train Semi-Truths
GenImage TPR 0.771 1.000 0.965
TNR 0.928 1.000 0.261
Accuracy 0.849 1.000 0.613
F1 Score 0.836 1.000 0.714
Semi-Truths TPR 0.909 0.830 1.000
TNR 0.962 1.000 1.000
Accuracy 0.920 0.865 1.000
F1 Score 0.947 0.907 1.000
Importance of dataset coverage.
These results highlight two key takeaways. First, a well-trained supervised classifier like SwinV2 can detect even high-fidelity synthetic images if it has seen samples from that generator. Second, existing detection models trained on older benchmarks fail to generalize to unseen, state-of-the-art generators. This highlights the necessity of updated and evolving detection datasets, such as OpenFake, to match the capabilities of modern generative models.
Robustness to compression artifacts.
Real images (sourced from LAION-400M in our dataset) are typically compressed and carry authentic JPEG artifacts and blur. In contrast, synthetic images are high-resolution and minimally compressed. This mismatch creates an obvious signal: detectors may rely on compression differences instead of true semantic features, thus failing on compressed fakes. To assess this vulnerability, we implemented a data augmentation pipeline to degrade synthetic images during training, mimicking the distribution of real images, as done in previous work (Corvi et al., 2023; Wang et al., 2021; Cheng et al., 2024). This includes random resizing, Gaussian blur, JPEG compression, and Gaussian noise. We then evaluated the SwinV2 model trained with these augmentations on a fully compressed test set and got an overall F1 score of 0.992. This demonstrates that the model remains highly accurate even when the compression signal is neutralized.
@@
Using the out-of-distribution generators in Table˜3 (blue rows), which were collected from public web sources rather than generated by us, the SwinV2 trained on OpenFake shows the strongest transfer while keeping a high true-negative rate on real images. The SwinV2 trained on GenImage is competitive on several open-source families but lags on newer proprietary models, and the Semi-Truths model attains high TPRs yet mislabels many real images, so its apparent OOD gains are not reliable. Cross-benchmark tests in Table˜4 reinforce this: when evaluated on GenImage, the OpenFake model substantially outperforms the Semi-Truths model (Accuracy 0.849 vs. 0.613; F1 0.836 vs. 0.714), and when evaluated on Semi-Truths it exceeds the GenImage model (Accuracy 0.920 vs. 0.865; F1 0.947 vs. 0.907), despite being out-of-domain in both cases. Overall, the most transferable detector is the one trained on the broad, up-to-date distribution provided by OpenFake.

### v2

@@ 3 Case Study: Real-World Misinformation and Human Limits
Social media platforms have become critical channels for political discourse, and consequently, for amplifying deepfake disinformation. This raises a key question: how are deepfakes actually deployed within political conversations? As part of a subsequent study, and to investigate how deepfakes are used in political conversations and to later evaluate our detector in a realistic setting, we collected images from X, Reddit, and Bluesky. The collection spans the period immediately before and during the 2025 Canadian federal election, enabling analysis of synthetic media in a high-stakes, time-sensitive information environment.
We manually examined over 2000 randomly sampled images collected during a 72-hour period to better understand the types of visuals circulating on social media, identifying 163 deepfakes (see Section˜6.1 for more details). Many of these prominently featured political figures. Fabricated scenarios involving leading candidates distorted public perception, and even when some deepfakes were clearly artificial (Figure˜2a), they reinforced existing biases more effectively than textual misinformation alone [22, 27, 71]. When photorealistic deepfakes aligned with viewers’ prior beliefs, the risk was higher, as illustrated in Figure˜2c. Beyond portraits, misinformation extended to political symbols, banners, and manipulated depictions of protests or disasters (Figure˜2d,e), often paired with misleading text (Figure˜2b). These findings inform the construction of our dataset: to support generalizable detection, a deepfake benchmark must move beyond faces to capture the full breadth of misleading visual content. While this dataset serves here to illustrate the variety of real-world deepfakes, in Section˜6.1 we also use it as a small in-the-wild evaluation set. Importantly, none of these images or their captions are included in training or generation, ensuring a clean separation between evaluation and benchmark design.
@@ 3.1 Human Perception Study
To assess the difficulty of detecting deepfakes generated by different models, we conducted a simple human study. A total of 100 participants completed the survey, each viewing a randomized set of 24 images. The set consisted of 12 real photographs and 12 synthetic images—2 from each of the six generation models: GPT Image 1 (OpenAI) [58], Imagen 3 (Google) [4], Grok 2 (xAI), Flux.1.0-dev (Black Forest Labs) [39], Stable Diffusion 2.1 [62], and Stable Diffusion 3.5. All synthetic images were generated from the same prompts as their real counterparts, using the text automatically extracted by our pipeline described in Section˜4. Each prompt was only shown once to a given participant. This ensured that responses reflected a fair and diverse exposure across the dataset. In total, the survey contained 168 unique images. The survey description, code, and images are available at https://github.com/vicliv/deepfake-survey.
Source
@@
Flux.1.0-dev 2024 Open 0.689 0.562 0.792 0.812 0.917 1.000 1.000
SDv3.5 2024 Open 0.709 0.521 1.000 0.938 0.750 1.000 1.000
SDv2.1 2022 Open 0.879 0.646 1.000 0.938 0.917 1.000 1.000
Grok 2 2024 Proprietary 0.811 0.771 1.000 0.938 0.208 1.000 1.000
Overall Accuracy 0.714 0.524 0.417 0.427 0.804 0.750 1.000
@@
Table 2: Survey results showing human true positive rate (TPR) for synthetic images and true negative rate (TNR) for real images. A score near 0.50 indicates chance-level performance. SwinV2 trained on OpenFake achieves high accuracy on its in-distribution generators.
@@
The results in Table˜2 highlight key insights into synthetic image realism and human perception. Imagen 3 from Google achieved the lowest human accuracy (48.5%), equivalent to random guessing. GPT Image 1, OpenAI’s recent model, was similarly deceptive, with nearly one-third of its images undetected. In contrast, Stable Diffusion 2.1 (SDv2.1, which is widely used in current benchmarks, had the highest detection rate (87.9%) and was easily flagged by all detectors. These findings suggest that while open-source models remain relatively easy to spot, advanced proprietary models demonstrate exceptional realism and consistency. This underscores the need to include such models in benchmarks, as current deepfake detection datasets trained on older or open-source models fail to capture the new quality standards set by proprietary systems.
@@
However, this perceptual difficulty is not limited to humans. Baseline deepfake detectors also failed to consistently identify these advanced fakes. The CLIP-D-10k+ method [17], which fine-tunes a linear classifier on CLIP embeddings, performed close to random on several models (e.g., 45.8% on Imagen 3, and GPT Image 1), and failed to distinguish real images entirely (47.9% TNR). Similarly, the Corvi2023 method [16], which uses curated handcrafted features for detection, fared better on some open-source models (e.g., 100% on SDv3.5), but completely failed on real images (0.0% TNR) and newer proprietary content like GPT Image 1 (50.0%). In contrast, our SwinV2 baseline, trained directly on a curated mix of real and synthetic images from models included in our dataset, achieved 100% overall accuracy and perfect performance on its in-distribution models. While this result demonstrates that deep networks can learn to detect even the most realistic fakes with sufficient supervision, it also highlights the brittleness of existing methods when faced with novel or unseen generative sources. The two additional SwinV2 baselines confirm this pattern: models trained on smaller or differently distributed data (GenImage and Semi-Truths) showed weaker generalization to modern deepfakes, with accuracy dropping to 80.4% and 75.0%, respectively. These results reinforce the importance of training on a large, diverse and higher-quality dataset like OpenFake to ensure robustness against emerging generation techniques.
@@ 4 Dataset Overview & Collection
OpenFake combines a large repository of real images with a diverse collection of high-quality synthetic counterparts generated by multiple state-of-the-art models. In Appendix A, Table˜5 presents in details the key statistics of the dataset. Figure˜4 offers a qualitative view of the underlying distribution of some of the images. The substantial overlap between real and synthetic samples in the CLIP feature space highlights their semantic alignment, suggesting that synthetic images effectively mimic the distribution of real-world content. This shows that the prompt generation pipeline is working as intended. Figure˜1 presents an overview of the data collection and generation process.
@@
We extract metadata from the LAION-400M dataset [66], which we selected due to its broad representation of internet-sourced images—the same domain where visual misinformation typically circulates. Additionally, this dataset was likely included in the training data of the text-to-image models used to generate the synthetic images, which should theoretically make it more difficult for detectors to distinguish between real and fake images. More importantly, these images preserve real-world compression artifacts, which are crucial for training detectors that operate in the wild. While LAION may contain some synthetic images, we expect this contamination to be minimal, as the dataset primarily consists of content from 2014–2021, before the public release of diffusion models in 2022. After scraping LAION, we filter image–caption pairs using a vision–language model (Qwen2.5‑VL [3]). As described by the prompt used to query the model in Appendix˜C, an image is retained if it is identified as depicting either (i) real human faces or (ii) politically salient or newsworthy events. For every retained image, we generate a more detailed caption to use as prompt input for text-to-image models. These $3\,\mathrm{M}$ prompts are also publicly released and form the basis of the prompts shown to users of our crowdsourcing platform OpenFake Arena.
We filtered real images using Qwen2.5-VL, selected for its trade-off between speed and quality. To prevent detection shortcuts, we excluded LAION images smaller than $\times512$ pixels from the released train/test sets, as lower resolutions introduced compression artifacts that made detection artificially easier. Full prompts used for filtering and captioning are provided in Appendix˜C, and additional details on generation and compute resources are in Appendix˜E.
@@
We generated images from a diverse set of state-of-the-art generators using samples from our prompt bank: Stable Diffusion 1.5/2.1/XL/3.5 [62], Flux 1.0-dev/1.1-Pro/Schnell [39], Midjourney v6/v7 [53, 54], DALL·E 3 [57], Imagen 3/4 [4, 26], GPT Image 1 [58], Ideogram 3.0 [33], Grok-2 [75], HiDream-I1 [32, 31], Recraft v3 [61], Chroma [49], and 10 community variants (Finetuned or LoRA) of Stable Diffusion 1.5/XL and Flux-dev. All images are produced at $\sim$1 MP resolution with varied aspect ratios (9:16, 16:9, 1:1, 2:3, 3:4, etc.), mirroring common social-media formats. Because several proprietary sources impose “non-compete” clauses, those subsets are released under a non-commercial license.
@@
We construct the train/test split by sampling 1,000 images per generative model (with the exception of out-of-distribution models, which contribute between 200 and 600 images) along with the corresponding number of real images, yielding a test set of roughly 60,000 images (about 3% of the dataset assuming balanced classes). The remaining images are allocated to the training set. To ensure balance, each model is equally represented in the test split, and real images are matched accordingly. The rest of the real images and prompts are provided in a CSV file, with the real images accessible through their URLs. As OpenFake Arena expands and more synthetic images are collected, additional real images will be incorporated to preserve parity between real and synthetic domains in the train/test splits. All assets are hosted on the HuggingFace Hub. All images have their associated prompts and model name as metadata, which can be used for model attribution.
@@ 5 Crowdsourced Adversarial Platform
The Arena features a real-time leaderboard to gamify the experience and incentivize participation. The detector is periodically retrained with newly collected data, enabling continual improvement. Submitted images are periodically reviewed. This human-in-the-loop setup transforms model drift from a challenge into a feature, allowing the benchmark to evolve organically alongside the state of generative models. Implementation details and screenshots of the arena are in Appendix˜F.
@@ 6 Baseline Detector Benchmarks
Our primary detector is SwinV2-Small [48], a hierarchical vision transformer that has achieved state-of-the-art results on large-scale classification tasks. We adopt it as the backbone for our supervised detector, trained on OpenFake. In addition, we evaluate two SwinV2 variants trained on external datasets (GenImage and Semi-Truths, chosen based on relevance from Table˜5), a ConvNeXt baseline trained on DRCT taken from [11], and an EfficientNet-B4 baseline trained on FaceForensics++ from [76]. For semi-supervised detection, we include the CLIP-Based Synthetic Image Detector [17], which applies a linear probe over CLIP embeddings with minimal training data. For zero-shot detection, we use InternVL [14] directly without finetuning. [^4] We use the InternVL3-38B; Code for all scripts and baselines https://github.com/vicliv/OpenFake Finally, we also test the handcrafted detector of Corvi et al. [16] and a hybrid fusion baseline that averages predictions from CLIP and Corvi2023. Together, these baselines cover a diverse range of architectures, training regimes, and prior benchmarks, allowing us to compare detectors trained on OpenFake against both legacy and contemporary approaches.
OpenFake SwinV2 GenImage SwinV2 S.-Truths SwinV2 DRCT ConvNeXt FF++ EffNet-B4 CLIP- D-10k+ DMD Corvi’23 InternVL-3 (zero-shot)
Real (TNR) 0.995 0.955 0.689 0.777 0.516 0.703 0.998 0.431
SD 1.5 1.000 0.936 1.000 0.447 0.529 0.579 0.000 0.849
SD 2.1 1.000 0.998 0.999 0.482 0.453 0.717 0.011 0.900
SD XL 1.000 0.956 1.000 0.426 0.507 0.438 0.001 0.814
SD 3.5 1.000 0.982 1.000 0.324 0.466 0.406 0.000 0.796
@@
Results on OpenFake.
As shown in Table˜3, the SwinV2 trained on OpenFake is near-perfect, confirming that modern classifiers are highly reliable when trained on the evaluation distribution. Separately, we evaluate robustness to compression artifacts and, with artifact-matching augmentation, the same model attains an F1 of 0.992 on a fully compressed test set (see Section˜B.2 for more details). The SwinV2 trained on GenImage is the next strongest and handles many shared open-source families, but degrades on newer proprietary generators (Grok 2, Midjourney 6, Flux-1.1 pro, DALL·E 3), reflecting a distribution gap in the fake images. The SwinV2 trained on Semi-Truths achieves high TPRs yet misclassifies many real images, which is consistent with its training data focused on edits rather than full generation. Legacy or narrow baselines (ConvNeXt/DRCT, EffNet-B4/FF++) underperform, CLIP-D-10k+ is middling, and Corvi2023 largely predicts “real,” while zero-shot InternVL is better than older baselines but still trails supervised models. Overall, robust performance requires training on the correct, broad, and up-to-date image distribution.
@@
Using the out-of-distribution generators in Table˜3 (blue rows), which were collected from public web sources rather than generated by us, the SwinV2 trained on OpenFake shows the strongest transfer while keeping a high true-negative rate on real images. The SwinV2 trained on GenImage is competitive on several open-source families but lags on newer proprietary models, and the Semi-Truths model attains high TPRs yet mislabels many real images, so its apparent OOD gains are not reliable. Cross-benchmark tests as seen in Table˜6 (Appendix B.1) reinforce this: when evaluated on GenImage, the OpenFake model substantially outperforms the Semi-Truths model (Accuracy 0.849 vs. 0.613; F1 0.836 vs. 0.714), and when evaluated on Semi-Truths it exceeds the GenImage model (Accuracy 0.920 vs. 0.865; F1 0.947 vs. 0.907), despite being out-of-domain in both cases. This suggests that while dataset coverage is the main driver of transferability, there is also some degree of cross-generator generalization, likely because different models share subtle artifacts; the broader the training distribution, the more likely a detector can recognize previously unseen generators.
6.1 Detector in the wild
Performance on benchmarks often differs from performance in real-world settings. For deepfake detection, evaluation in the wild is particularly challenging. Manually labelling real images is not always straightforward, since, as discussed in Section˜3.1, humans struggle to identify high-quality fakes. Real images are easier to validate because their authenticity can often be established through provenance cues such as credits from a reputable source, multiple photos of the same event, consistent backgrounds, or camera metadata. Images without any such evidence can be discarded as uncertain. The risk, however, is that the benchmark becomes trivial, with fakes limited to the easiest cases. With an experienced labeler, the aid of reverse image search, and contextual text (which may explicitly indicate AI generation), more difficult deepfakes can be identified. Using this approach, we constructed a small evaluation set of social media images, described in Section˜3, containing 1,057 real images and 163 labeled as deepfakes by us, and compared the performance of our SwinV2 baseline trained on OpenFake, GenImage, and Semi-Truths (Table˜4), since these were the only competitive baselines from Table˜3.
Metric Train OpenFake Train GenImage Train Semi-Truths
TNR 0.976 0.998 0.220
TPR 0.865 0.043 0.908
Accuracy 0.962 0.871 0.312
F1 Score 0.857 0.081 0.261
ROC–AUC 0.978 0.557 0.634
Table 4: Generalization of SwinV2 detectors trained on different benchmarks when evaluated on an in-the-wild social-media set (1,057 real, 163 fake; see Section˜3). Metrics include TNR (real) and TPR (fake). Training on OpenFake yields balanced performance, while GenImage and Semi-Truths show strong class biases.
The detector trained on OpenFake shows encouraging results. It produces very few false positives, meaning real images are rarely misclassified as deepfakes, while still identifying 86.5% of the fakes. Although the evaluation set is small and not without limitations—labels were verified, but some of the most difficult cases may have been discarded during curation (out of roughly 2000 candidate images, many were removed, including irrelevant real samples such as screenshots of text or drawings); the results strongly suggest that OpenFake offers superior real-world applicability compared to existing datasets. This is a promising outcome for the deepfake detection community. Expanding generator coverage and incorporating image edits, rather than only fully generated images, could further improve performance and move closer to practical, reliable detection systems.